OpenStudy (anonymous):

1. Tom and his best friend Edward are seeking to join a gym. Tom saw on television that Platinum Gym was having a grand opening in a few months. Tom is considering a membership at Platinum Gym. Platinum Gym charges a one-time registration fee of $60, and membership is $20 a month. Tom’s best friend Edward says he plans to join Super Fit because Super Fit only charges $10 a month but has a one-time registration fee of $150. (c) If Tom decides to join Platinum Gym and Edward joins Super Fit, who would pay less during the first month of membership? Who would pay less after a year of membership? S

OpenStudy (bibby):

how much would tom and edward pay for their first month of membership?

OpenStudy (anonymous):

Bibby im sorry!!! i really miss you!

OpenStudy (bibby):

cool

OpenStudy (anonymous):

please unblock me! i wanna be friends again! i dont know why i acted so stupid that day!!! im sorry!

OpenStudy (anonymous):

i love you

OpenStudy (bibby):

I'm just trying to help this dude with his math man, chill

OpenStudy (anonymous):

Tom must pay -y=20x+60 Edward must pay -y=10x=150

OpenStudy (anonymous):

wait no not negative

OpenStudy (bibby):

that's correct where x is the number of months, so where x=1, what are they paying?

OpenStudy (bibby):

oh I assumed you meant dashes

OpenStudy (anonymous):

yeah i did but then it dawned on me that it looked like a negative sign

OpenStudy (bibby):

silly you

OpenStudy (bibby):

so, about that first month?

OpenStudy (anonymous):

so for Edward y=80 and for tom im not quite sure

OpenStudy (bibby):

a month means x=1. so 10x+150=160

OpenStudy (anonymous):

okay

OpenStudy (bibby):

now after a year, tom pays 60+(12*20)=300 edwards pays 150+(12*10)=270

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry